OpenStudy (anonymous):

[sin(3a)]^2-[sin(a)]^2=sin(2a)sin(4a)

OpenStudy (anonymous):

\[\sin ^{2}\left(3 \alpha \right)-\sin ^{2}\left(\alpha \right)=\sin (2 \alpha) \sin (4 \alpha)\]